please answer quiclky



a) For the development of a given product, the cash flows can be attributed to multiple costs as well as some income sources. For the 3 period, the marketing cost is -124,730, the product revenues are 1,698,860, the production costs are -3,437,334. The discount rate is 10.


b)For the development of a given product, the cash flows can be attributed to multiple costs as well as some income sources. For the 8 year, the marketing cost is -264,687, the product revenues are 2,121,547, the production costs are -2,715,021. If the discounted rate is 21 percent per year, what is present value of this period’s cash flow?
Answer #1

(a)

Net cash flow in each year = Revenue - Marketing costs - Production costs = 1,698,860 - 124,730 - 3,437,334 = - 1,863,204

Present value of cash flows = - 1,863,204 x P/A(10%, 3) = - 1,863,204 x 2.4869

= - 4,633,602.03

(b)

Net cash flow (year 8) = Revenue - Marketing costs - Production costs = 2,121,547 - 264,687 - 2,715,021 = - 858,161

Present value of cash flow = - 858,161 x P/F(21%, 8) = - 858,161 x 0.2176

= - 186,760.84